(7%) Problem 13: A box slides down a plank of length d that makes an angle of θ with the horizontal as shown. μk is the kinetic coefficient of friction and us is the static coefficient of friction Otheexpertta.comm 33% Part (a) Enter an expression for the minimum angle (in degrees) the box will begin to slide. θmin-atan(με) Correct! 33% Part (b) Enter an expression for the nonconservative work done by kinetic friction as the block slides down the plank. Assume the box starts from rest and θ is large enough that it will move down the plank. wnc-μk m g cos(θ) d X Attempts Remain Feedback: is available 33% Part (c) For a plank of any length, at what angle θ (in degrees) will the final speed of the box at the bottom of the plank be 0.75 times the final speed of the box when there is no friction present? Assume μ,-0.36. Grade Summary Deductions .% Potential 100%
Answer #1

B)

d = distance traveled

Perpendicular to incline, force equation is given as

Fn = mg Cos\theta

kinetic frictional force is given as

fk = uk Fn

fk = ukmg Cos\theta

Work done by kinetic frictional force is given as

Wnc = fk d Cos180

Wnc = - ukmgd Cos\theta

c)

when there is friction :

hi = initial height from where the block starts = d Sin\theta

v = final speed at the bottom

using conservation of energy

mg hi - Wnc = (0.5) m v2

mg (d Sin\theta) - ukmgd Cos\theta = (0.5) m v2

2g (d Sin\theta) - 2ukgd Cos\theta = v2

v = sqrt(2g (d Sin\theta) - 2ukgd Cos\theta)                                             eq-1

When there is no friction :

v' = final speed at the bottom

using conservation of energy

mg hi = (0.5) m v'2

v' = sqrt(2g d Sin\theta)                                                     eq-2

it is given that

v = 0.75 v'

Using eq-1 and eq-2

sqrt(2g (d Sin\theta) - 2ukgd Cos\theta) = (0.75) sqrt(2g d Sin\theta)

sqrt(Sin\theta - uk Cos\theta) = (0.75) sqrt( Sin\theta)

sqrt(Sin\theta - (0.36) Cos\theta) = (0.75) sqrt( Sin\theta)

Sin\theta - (0.36) Cos\theta = (0.5625) Sin\theta

(0.4375) Sin\theta = (0.36) Cos\theta

tan\theta = 0.823

\theta = 39.5 deg